Marco invests his summer earnings of $2440. He invests part of the money at 8% per year and the rest at 7.5% per year. After one year these investments earned $193 in simple interest. How much did Marco invest at each rate?
My solution also has to have "let" statements.

Answer by mananth(16949)

Investment I 8.00% per annum ---x
Investment II 7.50% per annum ---y

x + y= 2440 ------------------------1
8.00% x + 7.50% y= = $193.00
Multiply by 100
8 x + 7.5 y= = $19,300.00 --------2

Multiply (1) by -8
we get

-8 x -8 y= = -19520.00

Add this to (2)

-0.5y= =-$220.00

divide by -0.5

y =$440.00 investment at 7.50%
Balance $2,000.00 investment at 8.00%
CHECK
$2,000.00 --------- 8.00% ------- $160.00
$440.00 ------- 7.50% ------- $33.00
Total -------------- $193.00

m.ananth@hotmail.ca